OpenStudy (anonymous):

multiply 9/7 times (3/5)

OpenStudy (anonymous):

27/35

OpenStudy (anonymous):

9/7 times (3/5)=9/7 *(3/5) 27/35

OpenStudy (anonymous):

to achieve the two anwser above you simply time the two top numbers toether(the numerators) and the bottom two numbers together(denominator), so thats 9*7 for the top+27 and 7*5=35

OpenStudy (anonymous):

sorry about the typing, keyboard is unresponsive